Showing all the steps, solve the following:
3x^3+x^2=4x

Can the below equation be solved? If yes, show how.
3x^2+4x=7

Answer by htmentor(1343)

Showing all the steps, solve the following:
3x^3+x^2=4x
=============================
3x^3+x^2=4x
This can be factored as
x(3x^2 + x - 4) = 0
So one of the solutions is x = 0
For the other two solutions we solve 3x^2 + x - 4 = 0
This can be factored as
(3x+4)(x-1) = 0
The solutions are x = -4/3, x = 1, and x = 0
==============================
Can the below equation be solved? If yes, show how.
3x^2+4x=7
=============================
3x^2 + 4x - 7 = 0
This can be factored as
(3x+7)(x-1) = 0
So the solutions are x=1, and x=-7/3